Assume today is December 31, 2013. Imagine Works Inc. just paid a dividend of $1.20 per share at the end of 2013. The dividend is expected to grow at 12% per year for 3 years, after which time it is expected to grow at a constant rate of 5.5% annually. The company's cost of equity (rs) is 9%. Using the dividend growth model (allowing for nonconstant growth), what should be the price of the company's stock today (December 31, 2013)? Round your answer to the nearest cent. Do not round intermediate calculations.

$
Answer #1

D1=(1.2*1.12)=1.344

D2=(1.344*1.12)=1.50528

D3=(1.50528*1.12)=1.6859136

Value after year 3=(D3*Growth rate)/(Required return-Growth rate)

=(1.6859136*1.055)/(0.09-0.055)

=50.8182528

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.344/1.09+1.50528/1.09^2+1.6859136/1.09^3+50.8182528/1.09^3

=$43.04(Approx).